# SoundMorpher: Perceptually-Uniform Sound Morphing with Diffusion Model

### v1

@@ 5.5 Discussion
Uninformative v.s. informative initial text prompt. Complex audio usually cannot easily yield precise information to users. For example, it is a challenge for non-professional users to describe the genre of a music. We conduct an ablation study for initial text prompt on music morphing to verify effectiveness of text-guided conditional embedding optimization. We use a general initial text prompt, ‘a sound clip of music composition.’, as an uninformative initial prompt. And we use the given text prompts in AudioLDM2  [^3] https://audioldm.github.io/audioldm2/ as informative inital prompts. As in Table 6, informative initial text prompts may help with resynthesis quality and further improves morph correspondence. Despite the improved resynthesis quality with informative initial text prompts, the results show a decline in morphing intermediateness and smoothness. One possible reason is the better resynthesis quality makes the resynthesis endpoints more distinct (i.e., larger semantic gap), which could lead to slight decline in intermediateness and smoothness. However, the performance difference on initial text prompts is not significant which illustrates effectiveness of conditional embedding optimization.
@@
Inference steps. In our experiment, we follow the configuration of Zhang et al. (2024) and set DDIM steps to 100. To verify whether DDIM steps affect SoundMorpher performances, we compare with 20 DDIM steps in Table 6. Larger inference step seems to help for reconstruction quality and slightly imporves morph quality, however, performance differences between inference steps are not significant. This indicates SoundMorpher is robust for inference steps, and we extend this ablation study on environmental sound morphing task in Appendix 11.1. Thus, we suggest selecting a suitable DDIM step to trade-off overall binary search algorithm time-consuming and morph quality.
Limitations. The current implementation of SoundMorpher based on AudioLDM2 with 16.0kHz sampling rate, which may limit output audio quality. The conditional embeddings optimization only applies to sounds that can be produced by AudioLDM2. Sound examples that close to white noise, such as pure wind blowing used in MorphGAN (Gupta et al., 2023) are not easily generated by AudioLDM2, which makes the conditional embedding optimization produce low quality resynthesis sounds. We also observed that input sounds with a large semantic gap (e.g., Church bells-Clock alarm in Table 2) result in lower morphing quality. Furthermore, we observed when two audios exhibit significant temporal structure differences, such as environmental sounds, SoundMorpher may produce abrupt transitions, see Appendix 11.4 and Figure 6 for further details.
6 Conclusion
@@
We propose SoundMorpher, a sound morphing method base on a pretrained diffusion model that produces perceptually uniform morphing trajectories. Unlike existing methods, we refined the sound morphing problem and explored an explicit connection between morph factor and perceptual stimuli of morphed results which offers better flexibility and higher morphing quality, making it adaptable to various morphing methods and real-world scenarios. We validate SoundMorpher by a series of objective quantitative metrics as well as mean opinion score study following criteria proposed by Caetano & Osaka (2012). These quantitative objective metrics may help to formalize future studies on sound morphing evaluation. Furthermore, we demonstrated that SoundMorpher can be applied to wide range of real-world applications in our experiments and conducted in-depth discussions. SoundMorpher also has the potential to achieve voice morphing, as its foundational model AudioLDM2 supports speech generation; however, we leave this exploration for future work.
Appendix
7 SoundMorper implementation details
This section provides further details on SoundMorpher pipeline and implementations.
7.1 Overall pipeline of SoundMorpher
The overall pseudo pipeline for SoundMorpher is provided in Algorithm 1, where the overall pipeline of SoundMorpher contains three parts: (1) Conditional embedding optimization. (2) Model adaptation. (3) perceptually uniform binary search with constant SPDP increment.
Algorithm 1 Pipeline of SoundMorpher
A pre-trained AudioLDM2 pipeline including a pre-trained VAE with a encoder $g_{\theta}$ and a decoder $g_{\phi}$, a pre-trained latent diffusion model $\epsilon_{\theta}$, a pre-trained T5 model $f_{\phi}$, and a pre-trained GPT-2 model $f_{\varphi}$. Learning rates $\eta_{1},\eta_{2}$. Source and target audios $x^{(0)}$ and $x^{(1)}$. An initial text prompts $y$. perceptually uniform interpolation number $N$. Tolerance error for binary search $\epsilon_{tolerance}$. Number of training steps for text inversion for conditional embedding optimization $T_{inv}$. Number of training steps for model adaptation $T_{adapt}$. LoRA rank $r$, Number of steps for DDIM $T$.
Start morph factor $\alpha_{start}=0$, end morph factor $\alpha_{end}=1$. Start perceptual point $p_{start}=[0,1]$, and end perceptual point $p_{end}=[1,0]$.
# Step 1: Text-guided conditional embedding optimization.
for $i$ from $1$ to $T_{inv}$ do
     Randomly sample time step $t$ and random noise $\epsilon_{t}\sim N(0,I)$.
     $E^{(0)}\leftarrow E^{(0)}-\eta_{1}\nabla_{E^{(0)}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta)$.
     $E^{(1)}\leftarrow E^{(1)}-\eta_{1}\nabla_{E^{(1)}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta)$.
end for
# Step 2: Model adaptation with LoRA.
for $i$ from $1$ to $T_{adapt}$ do
end for
# Step 3: Perceptual-uniform binary search with constant SPDP increment.
Obtaining initial latent states $z_{T}^{(0)}$ and $z_{T}^{(1)}$ by Equation 4
$p_{list}\leftarrow$ ConstantSPDP$(N,p_{start},p_{end})$triangleright Obtain target SPDP points by Algorithm 2
$\alpha_{list}\leftarrow\text{BinarySeach}(\alpha_{start},\alpha_{end},p_{list},\epsilon_{tolerance})$
for $\alpha$ in $\alpha_{list}$ do
     $E^{(\alpha)}\leftarrow(1-\alpha)E^{(0)}+\alpha E^{(1)}$
     for $t$ from T todo
     end for
end for
$x^{(\alpha)}\leftarrow\text{vocoder}(g_{\phi}(z_{0}^{(\alpha)}))$triangleright Decode latent variable and obtain audio waveform by a vocoder.
return $\{x^{(\alpha)}\}_{\alpha\in\alpha_{list}}$
7.2 Model adaptation with LoRA
In the task of image morphing, Yang et al. (2023) indicate adapting the model to the input pair helps to limit the degree of morphed variation by suppressing high-density regions that are not related to the given images. Compared to vanilla fine-tuning approaches, LoRA has advantages in training efficiency with injecting limited trainable parameters. The model adaptation can be defined as
$\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(0)},E^{(0)};\theta+\Delta\theta^{\prime})+\text{min}_{\Delta\theta^{\prime}}\mathcal{L}_{DPM}(z_{0}^{(1)},E^{(1)};\theta+\Delta\theta^{\prime})\text{ s.t. rank}(\Delta\theta^{\prime})=r$ (9)
where $r$ represents LoRA rank.
Unconditional Bias Correction. To achieve high text alignment during inference time, we use additional LoRA parameters $\Delta\theta_{0}$ with a small rank $r_{0}$ to perform bias correction, as
$\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(0)},\varnothing;\theta+\Delta\theta_{0})+\text{min}_{\Delta\theta_{0}}\mathcal{L}_{DPM}(z_{0}^{(1)},\varnothing;\theta+\Delta\theta_{0})\text{ s.t. rank}(\Delta\theta_{0})=r_{0}$ (10)
In our experiment, we set $r=4$ and $r_{0}=2$. Although Yang et al. (2023) provide a heuristic suggestion for setting the LoRA rank value for image morphing task, however, we further investigate the relationship between LoRA rank $r$ and method performance in Table 8 in sound morphing task and discussion in Appendix 11.2.
7.3 perceptually uniform binary search with constant SPDP increment
Algorithm 2 provides detail pesudocodes for how to implement perceptually uniform binary search with constant SPDP increment. This pseudo algorithm includes two steps, firstly, compute constant SPDP increment according to interpolte point number $N$ and obtain $N$ target SPDP points as $\{p_{i}\}_{i=1}^{N}$. Secondly, perform binary search to find correponding morph factor $\alpha$ series $\{\alpha_{i}\}_{i=1}^{N}$ according to $\{p_{i}\}_{i=1}^{N}$.
Algorithm 2 Pseudo algorithm for perceptually uniform binary search with constant SPDP increment
$\alpha_{start}$: starting alpha value; $\alpha_{end}$: ending alpha value; $N$: number of interpolations; source audio $x^{(0)}$; target audio $x^{(1)}$;
$p_{list}=[]$; $p_{start}=[0,1]^{T}$; $p_{end}=[1,0]^{T}$; $\alpha_{list}=[]$;
# Step 1: Obtain target SPDP points with constant increment.
Procedure ConstantSPDP$(N,p_{start},p_{end})$
for $i$ from $1$ to $N-1$ do $\triangleright$ Find target SPDP points with constant $\Delta p$.
     $t\leftarrow\frac{i}{N-1}$;
     $p_{i}\leftarrow(1-t)\times p_{start}+t\times p_{end}$
     $p_{list}\leftarrow p_{list}\cup[p_{i}]$;
end for
# Step 2: Perform binary search given target SPDP points with constant $\Delta p$.
Procedure BinarySearch($\alpha_{start},\alpha_{end},x^{(0)},x^{(1)},p_{list},\epsilon_{tolerance}$);
$\alpha_{list}\leftarrow[\alpha_{start}]$;
$\alpha_{cur}\leftarrow\alpha_{start}$;
for $p_{i}$ from $p_{1}$ to $p_{N-2}$ do
     $p_{target}\leftarrow p_{i}$
     $\alpha_{t1}\leftarrow\alpha_{cur}$;
     $\alpha_{t2}\leftarrow\alpha_{end}$;
     $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$;
     $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$triangleright Compute SPDP by Equation 7.
     while $|p_{mid}-p_{i}|>\epsilon_{tolerance}$ do $\triangleright$ Perform binary search
         if $p_{mid}>p_{target}$ then
              $\alpha_{t2}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         else
              $\alpha_{t1}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         end if
         $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
         $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
     end while
     $\alpha_{list}\leftarrow\alpha_{list}\cup[\alpha_{mid}]$triangleright Append the reult to $\alpha_{list}$.
end forreturn $\alpha_{list}$;
Algorithm 3 Pseudo algorithm for static morphing
Source audio $x^{(0)}$, target audio $x^{(1)}$, specific SPDP point $p$, tolerance error for binary search $\epsilon_{tol}$, number of steps for DDIM $T$, VAE decoder $g_{\phi}$.
$\alpha_{start}=0,\alpha_{end}=1$<

### v2

@@ V-E Discussion And Ablation Study
Uninformative v.s. informative initial text prompt. Complex audio usually cannot easily yield precise information to users. For example, it is a challenge for non-professional users to describe the genre of a music. We conduct an ablation study for initial text prompt on music morphing to verify effectiveness of text-guided conditional embedding optimization. We use a general initial text prompt, ‘a sound clip of music composition.’, as an uninformative initial prompt. And we use the given text prompts in AudioLDM2  [^6] https://audioldm.github.io/audioldm2/ as informative inital prompts. As in Table V, informative initial text prompts may help with resynthesis quality and further improves morph correspondence. Despite the improved resynthesis quality with informative initial text prompts, the results show a decline in morphing intermediateness and smoothness. One possible reason is the better resynthesis quality makes the resynthesis endpoints more distinct (i.e., larger semantic gap), which could lead to slight decline in intermediateness and smoothness. However, the performance difference on initial text prompts is not significant which illustrates effectiveness of conditional embedding optimization.
@@
Inference steps. In our experiment, we follow the configuration of [30] and set DDIM steps to 100. To verify whether DDIM steps affect SoundMorpher performances, we compare with 20 DDIM steps in Table V. Larger inference step seems to help for reconstruction quality and slightly imporves morph quality, however, performance differences between inference steps are not significant. This indicates SoundMorpher is robust for inference steps; therefore, we suggest selecting a suitable DDIM step to trade-off overall binary search algorithm time-consuming and morph quality.
Figure 7: Failure cases for SoundMorpher with $N=5$. The source (left first) and target (right first) sounds that have significant semantic difference in contents, this leads SoundMorpher to produce abrupt transitions.
TABLE VI: Ablation study on model adaptation with different LoRA rank $r$ on music morphing. Rank with -- represents results without LoRA model adaptation.
[…]
- 0.073 10.38 56.02 1.052 0.085 $\pm$ 0.054 0.198
4 0.047 9.85 56.25 0.847 0.068 $\pm$ 0.045 0.178
8 0.059 10.01 56.35 1.035 0.073 $\pm$ 0.051 0.180
16 0.059 9.95 56.14 1.058 0.075 $\pm$ 0.052 0.169
32 0.130 10.77 59.06 0.818 0.058 $\pm$ 0.082 0.158
Ablation study on LoRA rank. In this experiment, we conduct an ablation study on model adaptation with LoRA on the task of music morphing. We test SoundMorpher with different LoRA ranks as well as SoundMorpher without model adaptation. Following [52], we train LoRA parameters for 150 steps with 1e-3 learning rate. We also set unconditional bias correction with $r_{0}=2$ for 15 steps with 1e-3 learning rate. Table VI shows the results of SoundMorpher with different rank size on model adaptation settings. According to Table VI, SoundMorher without model adaptation has obvious performance drop on morphing correspondence compares results with LoRA model adaptation. Even though a higher LoRA rank has a slight improvement on perceptual reconstruction quality, however, SoundMorpher with $r=32$ indicates poor correspondence with large MFCCsE and large smoothness variance CDPAMstd. This result indicates that SoundMorpher with higher LoRA rank not lead to a better morphing quality. When $r=4$, SoundMorpher achieves the best performance on smoothness, and correspondence compared to $r=8$, $r=16$ and $r=32$. Therefore, we suggest LoRA rank for model adaptation in SoundMorpher shouldn’t be too large.
Limitations. Although SoundMorpher is the first open-world sound morphing method, offering high-quality morphed results and broad applicability, we have identified the following limitations: The current implementation of SoundMorpher based on AudioLDM2 with 16.0kHz sampling rate, which may limit output audio quality. The conditional embeddings optimization only applies to sounds that can be produced by AudioLDM2. Sound examples that close to white noise, such as pure wind blowing used in MorphGAN [12] are not easily generated by AudioLDM2, which makes the conditional embedding optimization produce low-quality resynthesis sounds. Although SoundMorpher produces high-quality sound morphing results, abrupt transitions can occur when the source and target sounds have significant temporal structure differences. A clear example of this is attempting to morph continuous environmental sounds with sounds that contain more silence, as in Fig. 7. Environmental sounds often consist of discrete and temporally separated events, such as a dog barking or a cat meowing, which have distinct and abrupt characteristics. These are inherently different from the more continuous and harmonically structured nature of music, where elements blend more fluidly over time. As a result, creating smooth transitions between such disjointed environmental sounds can be more challenging, leading to the perception of more abrupt or less natural transitions in the morphing process.
VI Conclusion
@@
We propose SoundMorpher, a open-world sound morphing method base on a pretrained diffusion model that produces perceptually uniform morph trajectories. Unlike existing methods, we refined the sound morphing problem and explored a more complex connection between morph factor and perception of morphed results which offers better flexibility and higher morphing quality, making it adaptable to various morphing methods and real-world scenarios. We validate SoundMorpher with a series of adapted objective quantitative metrics following three format sound morphing criterions proposed by [16]. These quantitative metrics may help to formalize future studies for direct comparison on sound morphing evaluation. Furthermore, we demonstrated that SoundMorpher can be applied to wide range of real-world applications in our experiments and conducted in-depth discussions. SoundMorpher also has the potential to achieve voice morphing, as its foundational model AudioLDM2 supports speech generation; however, we leave this exploration for future work.